Wave Equation: Expression

A taut clothesline has length L and a mass M. Atransverse pulse is produced by plucking one end of theclothesline. If the pulse makes nround trips along theclothesline in t seconds.(a) Find the expression for the speed of thepulse in terms of n, L, and t.
1Click here to preview your answer.Your answer is incorrect.

(b) Find the expression for the tension F in theclothesline in terms of the same variables and mass M.
2Click here to preview your answer.Your answer is incorrect.
Click here for help with symbolic formatting.
Answer #1
A taut clothesline has length L and a mass M. Atransverse pulse is produced by plucking one end of theclothesline. If the pulse makesn round trips along theclothesline in t seconds.(a) Find the expression for the speed of thepulse in terms of n, L, and t.

time period=n/t
frequency=1/T=t/n
so t/n=1/2L√T/m
now T=4L*L*m*t/n=4MLt/n

linear density=m=M/L
let the tension be T
speed=√T/m=√T*L/M=√4L*Lt/n=2L√t/n m/sec
1Click here to preview your answer.Your answer is incorrect.

(b) Find the expression for the tension F in theclothesline in terms of the same variables and mass M.
T=4L*L*m*t/n=4MLt/n N

2Click here to preview your answer.Your answer is incorrect.

answered by: daisie
Answer #2

a) Since the pulse makes n round trips, it travels 2L distance n times in time t.

∴the speed of the pulse is,

v = 2nL/t

b) Speed of the pulse is also given as,

v = √(FL/m)

∴ F = v2m/L

∴ F = (2nL/t)2(m/L) = (2n/t)2mL


answered by: nobody